John is 13 years less than 3 times as old as his son. Six years ago, he was 14 years more than twice as old as his son was then. How old is John?
Answer by conor.williams(1)

j + 13 = 3s
(j-6) - 14 = 2s
------------------------------
j = 3s - 13
j = (2s + 14) + 6
------------------------------
3s - 13 = 2s + 20
s = 33
------------------------------
j = 99 - 13
j = 86
------------------------------
so john is 86 and his son is 33